OpenStudy (jaweria):

Who ever can help me with my calculus questions, I will appreciate their help :)So can you help me?

OpenStudy (anonymous):

what is the question?

OpenStudy (jaweria):

r u still there @ satellite?

OpenStudy (anonymous):

Hey, what is the question?

OpenStudy (jaweria):

oh ok yeah the question is :

OpenStudy (jaweria):

\[f ' (x) = 2x ^{2/3}+4x ^{5/3}-3x+\sqrt{5}\]

OpenStudy (jaweria):

find f (x) when f ' (x) is given

OpenStudy (jaweria):

can you please help me to understand this whole question bcoz my professor didnt help tht much

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

you have to find the integral

OpenStudy (jaweria):

yes

OpenStudy (anonymous):

6/5x^(5/3)+3x^(8/3)-3/2x^(2)+sqrt(5)x+C

OpenStudy (anonymous):

ok, so it might help to have a table of integrals, but I can walk you through it the integral of x^n dx=(1/(n+1))*(x^(n+1))

OpenStudy (jaweria):

alright thanks :)

OpenStudy (anonymous):

so you use that for each of the pieces of f'(x), and you have to add a constant at the end so it is (6/5)*x^(5/3)+(3/2)*(x^8/3)-(3/2)*x^2-sq(5)*x+C

OpenStudy (anonymous):

right the 3 should be 3/2

OpenStudy (anonymous):

sometimes it is helpful to look at the before and after and figure out what you have to multiply by to get from the derivative to the original function

OpenStudy (jaweria):

ok

OpenStudy (jaweria):

so thats my answer?

OpenStudy (jaweria):

Anyone help !! I need to submit my homework today :(

OpenStudy (anonymous):

Before you take the integral you have to find the anti-derivative so if f'(x) is the derivative what is f(x). Then you can plug in the integral from a to b for x and solve from there. remember its F(b)-F(a)

OpenStudy (anonymous):

Does this help?